# The inverse Kalman filter

### v5

@@ 5 Concluding remarks
Several future directions are worth exploring. First, computing large matrix-vector products is ubiquitous, and the approach can be extended to different models, including some large neural network models. Second, it is an open topic to scalably compute the logarithm of the determinant of the covariance in (3) without approximation. Third, the new approach can be extended to estimate latent functions when forward equations are unavailable or computationally intensive in nonlinear or non-Gaussian dynamical systems, where ensemble Kalman filter [6, 34] and particle filter [21] are commonly used. Fourth, our approach can be extended for estimating and predicting multivariate time series [22] and generalized factor processes for categorical observations [4]. Finally, we aim to generalize the fast approach to scenarios with input dimensions larger than one.
@@ Appendix A. Kalman Filter
Let $\bm{\theta}_{t-1}\mid{\mathbf{y}}_{1:t-1}\sim\mathcal{MN}(\mathbf{m}_{t-1},\mathbf{C}_{t-1})$. Recursively for $t=2,\dots,{N}$:
@@
Denote $\mathbf{\tilde{y}}=(\tilde{y}_{1},\dots,\tilde{y}_{N})^{T}=\mathbf{L}^{-1}\mathbf{y}$, where $\mathbf{L}$ is the factor in Cholesky decomposition of a positive definite covariance ${\bm{\Sigma}}$, with ${\bm{\Sigma}}=\mbox{cov}[\mathbf{y}]$ and $\mathbf{y}=(y_{1},\dots,y_{N})^{T}$ defined as in (1). We have
$\tilde{y}_{t}=\frac{y_{t}-f_{t}}{{Q}^{\frac{1}{2}}_{t}},$ (A4)

### v6

@@ 5 Concluding remarks
Several future directions are worth exploring. First, computing large matrix-vector products is ubiquitous, and the approach can be extended to different models, including some large neural network models. Second, it is an open topic to scalably compute the logarithm of the determinant of the covariance in (3) without approximation. Third, the new approach can be extended to estimate latent functions when forward equations are unavailable or computationally intensive in nonlinear or non-Gaussian dynamical systems, where the ensemble Kalman filter [6, 34] and particle filter [21] are commonly used. Fourth, our approach can be extended for estimating and predicting multivariate time series [22] and generalized factor processes for categorical observations [4]. Finally, we aim to generalize the fast approach to scenarios with input dimensions larger than one.
@@ Appendix A. Kalman Filter
Let $\bm{\theta}_{t-1}\mid{\mathbf{y}}_{1:t-1}\sim\mathcal{MN}(\mathbf{m}_{t-1},\mathbf{C}_{t-1})$. Recursively for $t=2,\dots,{N}$ the following statements hold.
@@
Define $\mathbf{\tilde{y}}=(\tilde{y}_{1},\dots,\tilde{y}_{N})^{T}=\mathbf{L}^{-1}\mathbf{y}$, where $\mathbf{L}$ is the factor in Cholesky decomposition of a positive definite covariance $\bm{\Sigma}$}, with ${\bm{\Sigma}}=\mbox{cov}[\mathbf{y}]$ and $\mathbf{y}=(y_{1},\dots,y_{N})^{T}$ defined as in (1). We have
$\tilde{y}_{t}=\frac{y_{t}-f_{t}}{{Q}^{1/2}_{t}},$ (A4)